# Local Look-Ahead Guidance via Verifier-in-the-Loop for Automated Theorem Proving

### v1

@@ Appendix A Extended Discussions on Related Works
Another important step in theorem proving, and many other alignment related tasks, is to align the output of a given model based on the positive-negative preference pair of samples. There are online version of RLHF as in Bai et al. [2022], Ouyang et al. [2022], and offline versions with either an explicit reward model (e.g. [Christiano et al., 2017, Ziegler et al., 2019]) or an implicit reward model (e.g. DPO [Rafailov et al., 2024]) to encode the preference. The later method directly optimizes the model without training any independent reward model. In particular DPO uses Bradley-Terry preference model and the fact that the optimal solution to the KL-constrained reward maximization objective is known in closed form to simplify the training loss without dependence on a reward model [Rafailov et al., 2024]. Many follow-up works explored DPO variations and its shortcomings such as reward over-optimization [Gorbatovski et al., 2024, Ethayarajh et al., 2024].
Different forms of direct preference optimization have been explored in the recent literature for enhancing the reasoning capabilities of language models, in particular for mathematical problem solving, for example [Xiong et al., 2024, Yuan et al., 2024, Jiao et al., 2024, Cobbe et al., 2021, Lightman et al., 2023, Wang et al., 2024, Pang et al., 2024, Chen et al., 2024, Lu et al., 2024b]. The authors in Xiong et al. [2024] introduce a multi-turn version of DPO to use feedbacks from the verifiers, in their case the code interpreters, particularly for multi-turn scenarios, which requires trajectory-level preference optimization. The trajectory-level preference can be obtained by dataset labels of the gold answers or ORMs [Cobbe et al., 2021, Lightman et al., 2023]. A more fine-grained step-wise supervision can be used as in PRMs [Lightman et al., 2023] or by leveraging trajectory level preferences [Wang et al., 2024]. The idea of ORM and PRM has also been originally discussed in Uesato et al. [2022]. These works focus on mathematical reasoning, rather then formal theorem proving, which is the focus of our work. In our work, the theorem prover automatically provides the preference, which is used either during training or for an offline generated dataset. Furthermore, our work follows PRM philosophy as we consider feedback from Lean at each step of proof generation.
The authors in Jiao et al. [2024], Yuan et al. [2024] applied the original DPO or KTO by taking trajectory completion as a meta action. The online iterative versions of DPO originally designed for chat is adapted to achieve better CoT reasoning in Pang et al. [2024], Xie et al. [2024]. In the papers [Chen et al., 2024, Lai et al., 2024, Xie et al., 2024, Lu et al., 2024b], the authors have explored generating proxy step-wise labels for the intermediate steps of the reasoning trajectories.

### v2

@@ Appendix A Extended Discussions on Related Works
Another important step in theorem proving, and many other alignment related tasks, is to align the output of a given model based on the positive-negative preference pair of samples. There are online version of RLHF as in Bai et al. (2022); Ouyang et al. (2022), and offline versions with either an explicit reward model (e.g. Christiano et al. (2017); Ziegler et al. (2019)) or an implicit reward model (e.g. DPO Rafailov et al. (2024)) to encode the preference. The later method directly optimizes the model without training any independent reward model. In particular DPO uses Bradley-Terry preference model and the fact that the optimal solution to the KL-constrained reward maximization objective is known in closed form to simplify the training loss without dependence on a reward model Rafailov et al. (2024). Many follow-up works explored DPO variations and its shortcomings such as reward over-optimization Gorbatovski et al. (2024); Ethayarajh et al. (2024).
Different forms of direct preference optimization have been explored in the recent literature for enhancing the reasoning capabilities of language models, in particular for mathematical problem solving, for example Xiong et al. (2024); Yuan et al. (2024); Jiao et al. (2024); Cobbe et al. (2021); Lightman et al. (2023); Wang et al. (2024); Pang et al. (2024); Chen et al. (2024); Lu et al. (2024b). The authors in Xiong et al. (2024) introduce a multi-turn version of DPO to use feedbacks from the verifiers, in their case the code interpreters, particularly for multi-turn scenarios, which requires trajectory-level preference optimization. The trajectory-level preference can be obtained by dataset labels of the gold answers or ORMs Cobbe et al. (2021); Lightman et al. (2023). A more fine-grained step-wise supervision can be used as in PRMs Lightman et al. (2023) or by leveraging trajectory level preferences Wang et al. (2024). The idea of ORM and PRM has also been originally discussed in Uesato et al. (2022). These works focus on mathematical reasoning, rather then formal theorem proving, which is the focus of our work. In our work, the theorem prover automatically provides the preference, which is used either during training or for an offline generated dataset. Furthermore, our work follows PRM philosophy as we consider feedback from Lean at each step of proof generation.
The authors in Jiao et al. (2024); Yuan et al. (2024) applied the original DPO or KTO by taking trajectory completion as a meta action. The online iterative versions of DPO originally designed for chat is adapted to achieve better CoT reasoning in Pang et al. (2024); Xie et al. (2024). In the papers Chen et al. (2024); Lai et al. (2024); Xie et al. (2024); Lu et al. (2024b), the authors have explored generating proxy step-wise labels for the intermediate steps of the reasoning trajectories.
@@ Appendix B Additional Details
Figure 5: A simple proof example in Lean on the left shows the $min$ function is symmetric, and its proof tree is at the right. The numbers next to the boxes refer to the corresponding line numbers in the written proof.
@@
Appendix C MiniF2F and ProofNet Results
MiniF2F ProofNet
ReProver 26.5 13.8
ReProver* 37.7 15.3
LeanListener 36.9 14.4
[…]